the perimeter df a standard basketball court is 288ft.find the dimentions of the basketball court if it's length is 44ft.longer than it's width.
Answer by nerdybill(7384)

the perimeter df a standard basketball court is 288ft.find the dimentions of the basketball court if it's length is 44ft.longer than it's width.
.
Let w = width
then because "it's length is 44ft.longer than it's width"
w+44 = length
.
Perimeter = 2(width + length)
.
288 = 2(w + w+44)
288 = 2(2w+44)
288 = 4w+88
200 = 4w
50 feet = w (width)
.
length:
w+44 = 50+44 = 94 feet (length)
.
solution: dimensions are 50 feet by 94 feet